# Exploring the test of the no-hair theorem with LISA

### v1

@@ Abstract
In this study, we explore the possibility of testing the no-hair theorem with gravitational waves from massive black hole binaries in the frequency band of the Laser Interferometer Space Antenna (LISA). Based on its sensitivity, we consider LISA’s ability to detect possible deviations from general relativity (GR) in the ringdown. Two approaches are considered: an agnostic quasi-normal mode (QNM) analysis, and a method explicitly targeting the deviations from GR for given QNMs. Both approaches allow us to find fractional deviations from general relativity as estimated parameters or by comparing the mass and spin estimated from different QNMs. However, depending on whether we rely on the prior knowledge of the source parameters from a pre-merger or inspiral-merger-ringdown (IMR) analysis, the estimated deviations may vary. Under some assumptions, the second approach targeting fractional deviations from GR allows us to recover the injected values with high accuracy and precision. We obtain $(5\%,10\%)$ uncertainty on ($\delta\omega,\delta\tau)$ for the $(3,3,0)$ mode, and $(3\%,17\%)$ for the $(4,4,0)$ mode. As each approach constrains different features, we conclude that combining both methods would be necessary to perform a better test. In this analysis, we also forecast the precision of the estimated deviation parameters for sources throughout the mass and distance ranges observable by LISA.
@@ I Introduction
The first detection of gravitational waves (GWs) with LIGO [1, 2] produced by the coalescence of the black hole binary (BHB) GW150914 [3], marked the beginning of the GW astronomy era. At the same time, its detection opened a window to probe physics beyond the standard model and general relativity (GR). Since that first detection, the scientific community has been eager to test GR in the strong field regime [4, 5, 6, 7].
One of the most considered tests nowadays is the test of the no-hair theorem [8, 9] in the ringdown of BHB. In the last stage of the coalescence, once formed, the perturbed BH will settle down through the emission of GWs. In perturbation theory (PT), the radiation of a Kerr BH can be written as a superposition of damping sinusoidals [10, 11, 12, 13, 14, 15]. The strain in the plus and cross polarizations reads
@@
The equations describing a perturbed Kerr BH were first derived in [10]. A companion paper describes the subsequent radiation of gravitational waves [11] and introduces solutions in terms of spin-weighted spheroidal harmonics ${}_{s}S_{lmn}(a_{f}\tilde{\omega}_{lmn};\theta,\varphi)$ with $s=-2$ denoting the spin of the field. They depend on the phase $\varphi$ and on the angle $\theta$ between the normal of the orbital plane and the observer. Each harmonic is labeled by $(l,m,n)$ as polar and azimuthal angular numbers, and overtone respectively. Their eigenvalues are complex frequencies known as quasi-normal modes (QNMs) [14, 13]. Where the real part is the oscillation frequency, while the imaginary part corresponds to the inverse of the damping time
@@
The metric structure characterizes the QNMs’ values; thus, for a remnant Kerr BH, they depend only on the mass and spin ($M_{f},a_{f}$[10, 8]. In contrast, the amplitude and phase associated with each mode $(A_{lmn},\phi_{lmn})$ correspond to their excitation in the pre-merger phase, thus depending on the initial black holes’ parameters [16, 17].
@@
One approach to probe the no-hair conjecture is to use BH spectroscopy [14, 13, 15, 18, 19]. In GR, the values of the BH spectrum are defined solely by the mass and spin of the final BH. Therefore, when studying the spectrum of the remnant BH, one can trace back to those parameters [20, 21, 22, 23]. Then, the comparison of pairs of mass and spin obtained from different QNMs should be consistent with each other. In an alternative theory to GR, the values of the complex frequencies might deviate from those of GR [24, 25, 9]. Thus, the pairs of mass and spin derived from each QNM would not be consistent [16, 26]. Note that more than one QNM is required to perform this analysis. There exists another method where only one QNM is needed: it involves the comparison of estimated parameters in the pre-merger and post-merger phases.
@@
To this day, while the fourth observational run (O4) is ongoing, over a hundred sources have been detected by the LIGO-Virgo-Kagra (LVK) Collaboration [27, 28]. Nevertheless, the signature of QNMs seems to hide below the noise floor. Hints of spherical higher harmonics have been found [29] in the full inspiral-merger-ringdown (IMR) waveform for some events. Furthermore, the presence of the first overtone of the dominant QNM, i.e. $(2,2,1)$, has been inferred for the first event GW150914 [30, 31] (see the discussion on its detectability [32, 33, 34, 35]). As the sensitivity of current and future interferometers increases [2, 36, 37], we expect to detect more QNMs, hopefully already from the O4 analysis. However, the question of whether we can unmistakably observe a deviation from GR remains. Up to the O3 catalog, various analyses have been made with results always in agreement with GR [5, 4, 6, 7, 38, 39]. More reliable analyses to confidently discriminate any alternative theory would rely on a null hypothesis comparison in a Bayesian approach. However, this endeavor presents quite a challenge, since it would require that BH’s spectra be solved for alternative theories. Various developments in computing beyond-GR BH’s spectra have been undertaken, primarily for static or slowly rotating BHs in different alternative theories, e.g. [24, 25, 40, 41], see nonetheless [42] for rapidly rotating BHs in effective field theories. In the lack of waveform catalogs including deviations from GR, the best we can do is to allow for deviations of GR in a model-independent way, in the injection and search templates.
@@
In this exploratory study, we address the question of the extent to which LISA becomes distinctively sensitive to a deviation from GR in the ringdown phase of BHB coalescence. In a similar context, possible deviations from GR with LISA sources have been studied in [45] using the pSEOBNRv5HM waveform [46, 39]. In that work, the full IMR was used to find deviations present only in the ringdown. While using the full IMR is an advantage for low SNR sources, for high SNR, systematic errors in the full IMR waveform might bias the estimated parameters of the remnant if eccentricity, precession, or higher harmonics are not accounted for. In this work, we consider a more flexible prior knowledge of parameters, assuming a raw posterior distribution of the final BH parameters estimated from the IMR as uniform priors for this analysis. We use two approaches to assess LISA sensitivity to detect deviations from GR. We will also discuss the outcome of different assumptions on the priors.
The paper is divided as follows. In Sec. II, we describe the response of LISA to gravitational waves. Sec. III is dedicated to the methodology implemented, including a description of the data, the templates, and the likelihood computation. In Sec. IV we show and comment on the results. We then forecast how the GR test precision relates to SNR in Sec. V. We conclude in Sec. VI.
@@ II LISA response to waveforms
LISA consists of three spacecrafts (S/C) in heliocentric orbits and arranged in a triangular formation with arms of $2.5\times 10^{6}$ km. One reason for this particular setup is that with different combinations of individual phasemeter measurements, one can construct multiple synthetic interferometers.
@@
We use geometric units $c=G=1$ throughout the whole study. Lower indices $r$ and $s$ take values from 1 to 3 for the three S/C, representing the light-receiving and the light-sending spacecraft respectively. The S/C positions are defined by xr,s. $L_{rs}$ is the arm’s length between the two S/C. The vector k defines the direction of propagation of the GW, while $\hat{\textbf{n}}_{rs}$ denotes the direction of the beam. Lastly, $H_{rs}$ is the source’s gravitational strain projected on the arm. It reads
@@
where e+,× are the polarization tensors defined in the traceless-transverse gauge
@@
LISA does not work like a regular Michelson interferometer, as its interferometry is synthetically performed on the ground in the post-processing. Given that LISA arms will not have an equal length in space, nor would they be stationary, particular linear combinations with time delays of links are needed to cancel the noise produced from fluctuations in the laser [47, 48, 49, 50, 51, 52] among others noises. Those combinations are known as time delay interferometry (TDI) channels ($X$, $Y$, $Z$), and will be produced off-line, such that
@@
One should perform a cyclic permutation of the S/C indices to obtain channels $Y$ and $Z$. The S/C positions are required to compute the light travel time (LTT) of the armlength $L_{rs}(t)$ between them. For this reason, we use the orbits computed for each S/C. The LTT will affect the delay operators Eq. (10) as well as the projection of the strain onto the links Eq. (4).
To perform data analysis, optimal combinations of the channels $X$, $Y$ and $Z$ can be found to obtain quasi-orthogonal channels. They are defined as [47, 50, 48, 51, 49, 54]
@@
In this analysis, we work only with channels $A$ and $E$, as $T$ is almost blind to GWs [54].
@@ III Methodology
We aim to test for the presence of deviations from GR in the ringdown of MBHBs with a LISA prototype pipeline. To this end, we developed a code [55] capable of generating ringdown waveforms with the LISA response. Anticipating the detailed description that will be given in a separate paper [55], we describe the main features of the code in the following section.
@@ III.1 Data
The analysis procedure consists of generating a toy model describing the ringdown phase of a MBHB as the sum of damped oscillations with the response of LISA. The sum of damped oscillations in the ringdown of MBHB is given by Eq. (1), with
@@
The $\mathrm{GR}$ index indicates the values obtained within the GR framework. There are different ways to compute complex frequencies; we recommend [56] for a review on this topic. Here, we make use of the qnm package [57] which is based on a spectral eigenvalue approach [58].
Furthermore, $\Xi$ in the amplitude and phase stands for the intrinsic redshifted parameters of the progenitors $\Xi=(m_{1},m_{2},\chi_{1},\chi_{2})$, and $D_{l}$ is the luminosity distance of the source. The spheroidal harmonics in Eq. (1) can be decomposed as
@@
where we drop the dependence on $(a_{f}\tilde{\omega}_{lmn};\theta,\varphi)$ for clarity, and where
@@
The functions ${}_{s}Y_{lm}$ are the spin-weighted spherical harmonics and $d\Omega$ is the solid angle. It is easy to see in Eqs. (16) and (17) that one can recover the solution for the non-rotating (Schwarzchild) BH in the spherical harmonic basis when $a_{f}\rightarrow 0$.
Mode mixing arises naturally from the choice of representation in perturbation theory in terms of spheroidal harmonics, as opposed to the spherical harmonics which is the most natural representation in numerical relativity (NR). Various authors [59, 60] computed the values of spherical-spheroidal mixing coefficients, given by
@@
Since $\delta_{m^{\prime}m}$ is the Kronecker delta parameter, we can drop the prime in the first $m$. Thus, one can find these coefficients in the literature, written without the first $m$ at all $\sigma_{l^{\prime}lmn}$ or even written as $\mu_{ml^{\prime}ln}$. With this representation, the strain takes the form
@@
In our case, the amplitudes and phases used in Eq. (13) belong to fittings made by London et al. [61, 62], where the mode mixing is already accounted for. Thus we consider amplitudes labelled with $(l^{\prime}mlmn)$ as
@@
Therefore, in order to consider the following three QNMs, namely $[(2,2,0),(3,3,0),(4,4,0)]$, we include $[(2,2,2,2,0),(3,2,2,2,0),(3,3,3,3,0),(3,3,3,3,0),\\ (4,4,4,4,0)]$, see Eq. (21). Indeed, the resulting signal is a sum of decaying waves with amplitudes and phases for $lmn=[(2,2,0),(3,3,0),(4,4,0)]$.
We inject a fractional deviation of QNMs equal to $\delta\omega_{lmn}=[0.0,0.01,0.03]$ and $\delta\tau_{lmn}=[0.0,0.05,0.01]$ in the same QNM order, i.e. $lmn=[(2,2,0),(3,3,0),(4,4,0)]$. Of course, more QNMs could and should be added, but as a proof of concept, we decided to include only these three, leaving more complex searches for future work. We consider input data including a GW signal with and without noise. The sampling rate is set to 1 second as a compromise between the planned sampling rate of 0.25 s, the typical duration of the ringdown for a heavy source (about 7000 s for a mass of $10^{7}$ $M_{\odot}$) and the number of data points $8192$. The parameters used for the source injection are listed in Table 1. Note that we use $\iota$ as the inclination angle instead of $\theta$. We also write the ringdown SNR as well as the final parameters for the remnant BH, obtained with Eqs. (3.6) and (3.8) from [63].

### v2

@@ Abstract
In this study, we explore the possibility of testing the no-hair theorem with gravitational waves from massive black hole binaries in the Laser Interferometer Space Antenna (LISA) frequency band. Based on its sensitivity, we consider LISA’s ability to detect possible deviations from general relativity (GR) in the ringdown. Two approaches are considered: an agnostic quasi-normal mode (QNM) analysis and a method explicitly targeting the deviations from GR for given QNMs. Both approaches allow us to find fractional deviations from GR as estimated parameters or by comparing the mass and spin estimated from different QNMs. However, the estimated deviations may vary depending on whether we rely on prior knowledge of the source parameters from a pre-merger or inspiral-merger-ringdown analysis. Under some assumptions, the second approach targeting fractional deviations from GR allows us to recover the injected values with high accuracy and precision. We obtain $5\%,10\%$ uncertainty on $\delta\omega,\delta\tau$ respectively, for the $(3,3,0)$ mode, and $3\%,17\%$ for the $(4,4,0)$ mode. As each approach constrains different features, we conclude combining both methods would be necessary to perform a better test. In this analysis, we also forecast the precision of the estimated deviation parameters for sources throughout the mass and distance ranges observable by LISA.
@@ I Introduction
The first detection of gravitational waves (GWs) by the LIGO-Virgo-KAGRA (LVK) Collaboration [1, 2, 3, 4], produced by the coalescence of the black hole binary (BHB) GW150914 [5], marked the beginning of the GW astronomy era. At the same time, its detection opened a window to probe physics beyond the standard model and general relativity (GR). Since that first detection, the scientific community has been eager to test GR in the strong field regime [6, 7, 8, 9].
Indeed, GR can be tested, for instance, through the no-hair theorem by analyzing the ringdown phase. The no-hair theorem is a prediction of GR that states that any Kerr black hole (BH) is characterized by only two parameters: its mass and its spin ($M_{f},a_{f}$[10, 11, 12, 13]. In the last stage of the coalescence, i.e., the ringdown, the perturbed remnant BH will settle down through the emission of GWs. In perturbation theory (PT), the gravitational radiation of a Kerr BH can be written as a superposition of damped sinusoids [14, 15, 16, 17, 18, 19]. The GW strain in the plus and cross polarizations reads
@@
The equations describing a perturbed Kerr BH were first derived by Teukolsky in 1973 [14]. In collaboration with Press [15], he described the subsequent radiation of GWs, decomposing the solutions in terms of spin-weighted spheroidal harmonics ${}_{s}S_{lmn}(a_{f}\tilde{\omega}_{lmn};\theta,\varphi)$, with $s=-2$ denoting the spin of the field. These spheroidal harmonics depend on the eigenvalues $\tilde{\omega}_{lmn}$, the spin, the phase $\varphi$ and the angle $\theta$ between the normal of the orbital plane and the observer. Each harmonic is then labeled by $(l,m,n)$ as polar and azimuthal angular numbers and overtone, respectively. These numbers are integers taking values: $|s|\leq l,|m|\leq l,0\leq n$, where the (2,2,0) is, in general, the dominant carrier of energy. The eigenvalues of the Teukolsky master equation are complex frequencies $\tilde{\omega}_{lmn}$ known as quasi-normal modes (QNMs) [18, 17]. The real part of these solutions denotes the oscillation frequency, while the imaginary part corresponds to the inverse of the damping time:
@@
The QNMs’ values are characterized by the metric structure; thus, for a remnant Kerr BH, they depend only on the mass and spin [14, 15, 20, 21]. In contrast, the amplitude and phase associated with each mode $(A_{lmn},\phi_{lmn})$ are related to their excitation in the pre-merger phase, thus depending on the initial BHs’ parameters [22, 23].
@@
One approach to probe the no-hair theorem is to use BH spectroscopy [18, 17, 19, 24, 25]. In GR, the values of the BH spectrum, i.e. the QNMs, are defined solely by the mass and spin of the final BH. Therefore, when studying the spectrum of the remnant BH, one can recover those two parameters with a parametrization [20, 21, 26, 27]. Then, two pairs of mass and spin obtained from different QNMs should be consistent with each other. In an alternative theory to GR, the values of the complex frequencies might deviate from those of GR [28, 29, 30]. Thus, the pairs of mass and spin derived from each QNM in an alternative theory, might not longer be consistent with each other [22, 31]. Note that more than one QNM is required to perform this analysis. There also exists another method where only one QNM is needed; it involves the comparison of the estimated parameters of mass and spin in the pre-merger phase with the estimated parameters of mass and spin from the QNM [7, 6, 8, 9, 32].
@@
To this day, while the fourth observational run (O4) is ongoing, almost a hundred sources have been confidently observed by the LVK Collaboration [33], while another hundred sources are additionally potential candidates [34]. In several events, the characteristic ringdown waveform has been observed [9]. Moreover, small hints of spherical harmonics beyond the dominant ($2,2$) mode have been found for some events in [35], using the tidal effective one-body post-merger (TEOBPM) [36] waveform model. However, the signature of QNMs seems to hide below the noise floor, despite the possible presence of the first overtone of the dominant QNM, i.e. $(2,2,1)$, which has been inferred with low significance for the first event GW150914 [32, 37] (see the discussion on its detectability [38, 39, 40, 41], and further analyses [42, 43, 44, 45]). As the sensitivity of current and future interferometers increases [2, 46, 47], we expect to detect more QNMs, hopefully already from the O4 analysis. However, the question of whether we can unmistakably observe a deviation from GR remains. Up to the O3 catalog, various analyses have been made with results always in agreement with GR [7, 6, 8, 9, 48, 49]. More reliable analyses to confidently discriminate an alternative theory would rely on a null hypothesis comparison in a Bayesian approach. However, this endeavor presents quite a challenge since it would require that BH’s spectra should be solved for alternative theories. Various developments in computing beyond-GR BH’s spectra have been undertaken, primarily for static or slowly rotating BHs in different alternative theories, e.g. [28, 29, 50, 51], see nonetheless [52] for rapidly rotating BHs in effective field theories and [53, 54] for a new numerical approach to compute the QNM values in beyond-GR theories. In the lack of waveform catalogs including deviations from GR, the best we can do is to allow for deviations of GR in a model-independent way, both in the injection and search templates.
@@
In this exploratory study, we address the question of the extent to which LISA becomes distinctively sensitive to a deviation from GR in the ringdown phase of BHB coalescence. In a similar context, possible deviations from GR with LISA sources have been studied in [57] using the pSEOBNRv5HM waveform [58, 49]. In that work, the full inspiral-merger-ringdown (IMR) was used to find deviations present only in the ringdown. While using the full IMR is an advantage for low SNR sources, for high SNR, systematic errors in the full IMR waveform might bias the estimated parameters of the remnant if eccentricity, precession, or higher harmonics are not accounted for. In this work, we consider a more flexible prior knowledge of parameters, assuming a raw posterior distribution of the final BH parameters estimated from the IMR as uniform priors for this analysis. We use two approaches to assess LISA sensitivity to detect deviations from GR, namely an agnostic approach and a deviations approach targeting the deviations in specific QNMs. We will also discuss the outcome of different assumptions on the priors for the first case.
The paper is divided as follows. In Sec. II, we describe the response of LISA to GWs. Sec. III is dedicated to the methodology implemented, including a description of the data, the templates, and the likelihood computation. In Sec. IV, we show and comment on the results. We then forecast how the GR test precision relates to SNR in Sec. V. We conclude in Sec. VI.
@@ II LISA response to waveforms
LISA consists of three spacecrafts (S/C) in heliocentric orbits and arranged in a triangular formation with arms of $2.5\times 10^{6}$ km. One reason for this particular setup is that with different combinations of individual phasemeter measurements, one can construct multiple synthetic interferometers [59].
@@
We use geometric units $c=G=1$ throughout the whole study. Lower indices $r$ and $s$ take values from 1 to 3 for the three S/C, representing the light-receiving and the light-sending spacecraft, respectively. The S/C positions are defined by xr,s. $L_{rs}$ is the arm’s length between the two S/C. The vector k defines the direction of propagation of the GW, while $\hat{\textbf{n}}_{rs}$ denotes the direction of the beam. Lastly, $H_{rs}$ is the source’s gravitational strain projected on the arm. It reads
@@
where $\psi$ is the polarization angle, defined as the rotation angle along the line of sight between the source frame and the observational frame, set to zero in the following. The polarization tensors, e+,×, are defined in the traceless-transverse gauge
@@
LISA does not work like a regular Michelson interferometer, as its interferometry is synthetically performed on the ground in the post-processing. Given that LISA arms will not have an equal length in space, nor would they be stationary, particular linear combinations with time delays of links are needed to cancel the noise produced from fluctuations in the laser [60, 61, 62, 63, 64, 59] among others noises. The combinations that allow for the laser noise suppression are known as time delay interferometry (TDI) channels ($X$, $Y$, $Z$) and will be produced off-line, such that
@@
One should perform a cyclic permutation of the S/C indices to obtain channels $Y$ and $Z$. The S/C positions are required to compute the light travel time (LTT) of the armlength $L_{rs}(t)$ between them. For this reason, we use the orbits computed for each S/C from ESA LISA science orbit files [67] generated with the LISA orbits package [68]. The LTT will affect the delay operators Eq. (10) as well as the projection of the strain onto the links in Eq. (4).
To perform data analysis, optimal combinations of the channels $X$, $Y$, and $Z$ can be found to obtain quasi-orthogonal channels. They are defined as [60, 63, 61, 64, 62, 69]
@@
In this analysis, we work only with channels $A$ and $E$, as $T$ is almost blind to GWs in the low-frequency regime [69]. Even though that channel could provide complementary information about the source, the computational effort is unrewarding in this case.
@@ III Methodology
We aim to test for the presence of deviations from GR in the ringdown of MBHBs with a LISA prototype pipeline. To this end, we developed a code [70] capable of generating ringdown waveforms with the response of LISA. Anticipating the detailed description that will be given in a separate paper [70], we describe the main features of the code in the following section.
@@ III.1 Data
The analysis procedure consists of generating a toy model describing the ringdown phase of an MBHB as the sum of damped oscillations with the response of LISA. The sum of damped oscillations in the ringdown of MBHB is given by Eq. (1), with
@@
The $\mathrm{GR}$ index indicates the values obtained within the GR framework. There are different ways to compute complex frequencies; we recommend [71] and [72] for reviews on this topic. Here, we make use of the qnm package [73], which is based on a spectral eigenvalue approach [74].
Furthermore, the letter $\Xi$ in the amplitude and phase stands for the intrinsic redshifted mass and spin parameters of the progenitors $\Xi=(m_{1},m_{2},\chi_{1},\chi_{2})$, and $D_{l}$ is the luminosity distance of the source. The spheroidal harmonics in Eq. (1) can be decomposed as [15]
@@
where we drop the dependence on $(a_{f}\tilde{\omega}_{lmn};\theta,\varphi)$ for clarity, and where [15]
@@
The functions ${}_{s}Y_{lm}$ are the spin-weighted spherical harmonics, and $d\Omega$ is the solid angle. It is easy to see in Eqs. (16) and (17) that one can recover the solution for the non-rotating (Schwarzchild) BH in the spherical harmonic basis when $a_{f}\rightarrow 0$.
Mode mixing arises naturally from the choice of representation in PT in terms of spheroidal harmonics, as opposed to the spherical harmonics, which is the most natural representation in numerical relativity (NR). Various authors [75, 76] computed the values of spherical-spheroidal mixing coefficients, given by
@@
Since $\delta_{m^{\prime}m}$ is the Kronecker delta parameter, one can drop the prime in the first $m$. Thus, one can find these coefficients in the literature, written without the first $m$ at all $\sigma_{l^{\prime}lmn}$ or even written as $\mu_{ml^{\prime}ln}$. With this representation, the strain takes the form
@@
In our case, the amplitudes and phases used in Eq. (13) belong to parametrizations made by London et al. [77, 78], where the mode mixing is already accounted for. Thus, we consider amplitudes labeled with $(l^{\prime}mlmn)$ as
@@
Therefore, to consider the following three QNMs, namely $[(2,2,0),(3,3,0),(4,4,0)]$, we include $[(2,2,2,2,0),(3,2,2,2,0),(3,3,3,3,0),(3,3,3,3,0),\\ (4,4,4,4,0)]$, see Eq. (21). Indeed, the resulting signal is a sum of decaying waves with amplitudes and phases for $lmn=[(2,2,0),(3,3,0),(4,4,0)]$.
In consistency with current constrains on the fractional deviations [9] and the possible error on those parameters predicted for LISA [57] for a source with an SNR $\leq 600$, we inject fractional deviations of QNMs equal to $\delta\omega_{lmn}=[0.0,0.01,0.03]$ and $\delta\tau_{lmn}=[0.0,0.05,0.1]$ in the following order $lmn=[(2,2,0),(3,3,0),(4,4,0)]$. Of course, more QNMs could and should be added, but as a proof of concept, we decided to include only these three, leaving more complex searches for future work.
We consider input data including a GW signal with and without noise. The sampling rate is set to 1 second as a compromise between the planned sampling rate of 0.25 s, the typical duration of the ringdown for a heavy source (about 7000 s for a mass of $10^{7}$ $M_{\odot}$) and the number of data points $8192$. The parameters used for the source injection are listed in Table 1. Note that we use $\iota$ as the inclination angle instead of $\theta$. We also write the ringdown SNR as well as the final parameters for the remnant BH, obtained with Eqs. (3.6) and (3.8) from [79].